Finding the real solutions to the equation:
Please help me solve this: 5x^2/3-24x^1/3-5=0
I got to the second part of this answer which is 125...but I cant figure out the fraction part. The answer should look something like this: -number/number,125

Found 2 solutions by Alan3354, lynnlo:
Answer by Alan3354(69443)

Finding the real solutions to the equation:
Please help me solve this: 5x^2/3-24x^1/3-5=0
-------------------
Sub u for x^(1/3)
------
5u^2 - 24u - 5 = 0
(5u + 1)*(u - 5) = 0
u = 5
x^(1/3) = 5
x = 125
--------------
u = -1/5
x = -1/125

Answer by lynnlo(4176)

5x^2/3-24x^1/3-5=0
5x^2/3-8x^1/1-5=0
5x^2/3-8x^1-5=0
5x^2-8x-5=0
x=12+square root219/5 or12-square root219/5
x~5.3597,,,,,,,,,,,,-0.5597